OpenStudy (anonymous):

if you deposit $1000 in to an account that pays 4% interest compounded continuously, how long will it take the account to grow to $2000

OpenStudy (anonymous):

The formula for interest compounded continuously is A= Pe^(rt). Substitute in your numbers, solve for t.

OpenStudy (anonymous):

2000 = 1000 (1+ 4/100)^n (1.04)^n = 2 ??

OpenStudy (anonymous):

"2000 = 1000 (1+ 4/100)^n" Where did you get that? Lol you leave e as e. There should be a button on your calculator that is labeled "e'" 2000=1000e^(0.04t)

OpenStudy (anonymous):

ooh i see